OpenStudy (anonymous):

Given the two points (0,4) and (2, 6), what would the numerator of the equation for slope be? 6 - 4 6 - 2 2 - 4 2 - 0

OpenStudy (anonymous):

what r we supposed to do here

OpenStudy (anonymous):

thats my qoustion lol

OpenStudy (anonymous):

Find the numerator

OpenStudy (anonymous):

numerator of.......?

OpenStudy (anonymous):

Y2-y1/x2-x1

OpenStudy (anonymous):

so whats the aanswer?

OpenStudy (anonymous):

no idea and i dont want to give u a false answer

OpenStudy (anonymous):

6-4/2-0 : 2/2

OpenStudy (the_fizicx99):

The slope formula is: M = (Y2-Y1) / (X2-X1) M = 6-4/2-0 M = 2/2 M = 1

OpenStudy (anonymous):

Slope: 2/2 or 1

OpenStudy (the_fizicx99):

It's A *

OpenStudy (anonymous):

4-6 because your y1 is 4 and your y2 would be 6

OpenStudy (the_fizicx99):

Actually that would only work if its y1-y2/x1-x2 But by looking at the answers they wanted y2-y1/x2-x1

OpenStudy (anonymous):

Its A